OpenStudy (amrit):

Evaluate: 100(1 + 0.08)

OpenStudy (anonymous):

= 100 + 100* 0.08

OpenStudy (anonymous):

how would you round that to the nearest tenth

OpenStudy (anonymous):

the answer is 100 + 8 = 108

OpenStudy (anonymous):

\[100+8=108\] rounded to nearest tenth is 110

OpenStudy (anonymous):

Evaluate: 100(1 - 0.02) to the 10 power

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Evaluate: 100(1 - 0.02)

OpenStudy (anonymous):

Tell whether the function shows growth or decay. Then graph. f(x) = 32(0.5x)